# Towards Safer Chatbots: A Framework for Policy Compliance Evaluation of Custom GPTs

### v1

@@ Abstract
Large Language Models (LLMs) have gained unprecedented prominence, achieving widespread adoption across diverse domains and integrating deeply into society. The capability to fine-tune general-purpose LLMs, such as Generative Pre-trained Transformers (GPT), for specific tasks, domains, or requirements has facilitated the emergence of numerous Custom GPTs. These tailored models are increasingly made available through dedicated marketplaces, such as OpenAI’s GPT Store. However, the black-box nature of the models introduces significant safety and compliance risks. In this work, we present a scalable framework for the automated evaluation of Custom GPTs against OpenAI’s usage policies, which define the permissible behaviors of these systems. Our framework integrates three core components: (1) automated discovery and data collection of models from the GPT store, (2) a red-teaming prompt generator tailored to specific policy categories and the characteristics of each target GPT, and (3) an LLM-as-a-judge technique to systematically analyze each prompt-response pair for potential policy violations.
We validate our framework with a manually annotated ground truth, and evaluate it through a large-scale study evaluating 782 Custom GPTs across three categories: Romantic, Cybersecurity, and Academic GPTs. Our manual annotation process achieved an F1 score of 0.975 in identifying policy violations, confirming the reliability of the framework’s assessments. The evaluation results reveal that 58.7% of the analyzed models exhibit indications of non-compliance, exposing weaknesses in the GPT store’s review and approval processes. Furthermore, our findings indicate that a model’s popularity does not correlate with its compliance level, and non-compliance issues largely stem from behaviors inherited from base models rather than user-driven customizations.
Our framework provides a viable solution for large-scale policy compliance evaluation, enhancing the safety of personalized LLMs and supporting systematic enforcement of usage policies. We believe this approach is extendable to other chatbot platforms and policy domains, improving LLM-based systems safety.
@@ 1 Introduction
The advent of Transformer architecture and its self-attention mechanism [1] marked a turning point in the field of Natural Language Processing (NLP), enabling the development of Large Language Models (LLMs). These models revolutionized NLP by introducing parallelized processing, significantly improving training speed and scalability compared to prior architectures such as Recurrent Neural Networks (RNNs) and Long Short-Term Memory networks (LSTMs).
Building on this foundation, OpenAI developed the Generative Pre-trained Transformer (GPT) series, combining unsupervised pre-training with supervised fine-tuning to create GPT-1 [2]. Subsequent iterations, including GPT-2 [3] and GPT-3 [4], demonstrated remarkable advancements, with each version achieving greater performance through larger training datasets and architectural refinements. However, it was the release of ChatGPT in 2022, integrating an optimized GPT-3.5 model for dialogue, that marked a paradigm shift in the adoption of LLMs. ChatGPT showcased the versatility of these models across diverse applications, including code generation, education, marketing, and task automation. Its rapid adoption highlights the widespread appeal and utility of this technology, becoming the fastest platform to reach 100 million users within just two months of its launch [5].
AI agents emerged shortly thereafter as a strategic objective for the industry, aiming to develop LLMs specialized in specific tasks and optimized for targeted performance. This approach quickly gained traction among major technology companies, which recognized the potential of these agents to seamlessly integrate into workflows such as software development and customer support [6, 7]. The development of these agents was primarily enabled by fine-tuning, a technique that customizes base models by retraining them on task-specific datasets, enabling tailored behavior and improved performance in specialized contexts.
Building on the maturation of fine-tuning techniques, OpenAI introduced GPTs [8]—a.k.a. Custom GPTs—in November 2023, providing an accessible platform for users to personalize LLMs. By providing a user-friendly interface, OpenAI lowered technical barriers, enabling paid users to customize base models according to their specific requirements. To further enhance the accessibility and utility of these models, OpenAI launched the GPT store, a centralized marketplace where users can publish and share their Custom GPTs with a global audience. However, prior to publication, these models must pass a review process—both automated and manual—designed to ensure their compliance with OpenAI’s usage policies and to mitigate potential safety risks [9].
As the adoption of Custom GPTs expanded, OpenAI implemented usage policies [9] to define the boundaries of permissible behavior for Custom GPTs and mitigate risks associated with user-generated models. These policies explicitly prohibit GPTs from engaging in tasks such as providing legal or medical advice, generating malware, or performing other harmful activities. Despite these safeguards, Custom GPTs that contravene these policies seem to be present in the platform. For instance, Custom GPTs designed for romantic interactions—expressly prohibited under these policies—have proliferated within the platform (see Figure 1). This phenomenon underscores significant shortcomings in the review and approval processes, highlighting the challenges of scaling oversight mechanisms for user-generated content. Given the sheer volume of Custom GPTs in the GPT Store, manual efforts alone are insufficient to address the scale of the problem effectively. The challenges of scalability and consistency in enforcing compliance demand an automated solution capable of systematically evaluating models.
@@
Research Questions and Contributions
To address this gap and enhance the safety of the GPT store, our work is guided by the following research questions:
@@
  • RQ3: How does the customization of GPTs impact their compliance with usage policies, compared with the baseline models?
@@
This study addresses the outlined research questions through the following contributions. First, we propose and publicly release a novel and scalable framework for the automated evaluation of Custom GPTs’ compliance with OpenAI’s usage policies. The framework integrates multiple modules, including a custom GPT discovery tool, a tailored red-teaming prompt generator, and an LLM-as-a-judge assessment mechanism, enabling end-to-end compliance evaluation, as detailed in §2. Second, we validate the framework using a manually crafted ground truth dataset and demonstrate its effectiveness through a large-scale study involving 782 Custom GPTs from OpenAI’s GPT store, uncovering patterns of compliance and non-compliance across key thematic categories (Romantic, Cybersecurity, and Academic GPTs), as presented in §3 and §4, respectively. Third, we provide a statistical analysis of compliance results, offering insights into factors such as thematic focus and GPT popularity, and further analyze specific cases of non-compliance, highlighting their broader implications (§5).
@@
We detail the process of disclosing non-compliant Custom GPTs to OpenAI in §6 and discuss our findings’ implications in §7. To contextualize our contributions and situate our work within the broader research landscape, we discuss related studies on policy compliance evaluation and adversarial testing in §9. Finally, we conclude with a summary of contributions and directions for future research in §10.
@@ 2.1 Overview
This section provides an overview of the proposed framework designed for the systematic evaluation of Custom GPTs’ compliance with usage policies. Figure 2 presents a high-level architecture diagram outlining the framework’s primary modules and workflow, encompassing the entire evaluation lifecycle—from the identification of Custom GPTs to their compliance assessment.
Figure 2: Framework Architecture for Automated Policy Compliance Assessment of Custom GPTs. The framework comprises interconnected modules designed to evaluate Custom GPTs’ compliance with usage policies. In Phase I (Stage 1), the Custom GPT Interactor module retrieves a list of Custom GPTs, along with their associated metadata. The Red-Teaming Prompts Generator then generates tailored prompts to test each GPT against specific policy criteria (Stage 2). In Phase II, these prompts are transmitted to the Custom GPTs, and their responses are collected (Stage 3). The Compliance Assessment module evaluates the generated conversation transcripts using the LLM-as-a-judge technique (Stage 4), providing fine-grained compliance results. Finally, the Orchestrator stores and logs the compliance evaluation results for subsequent analysis (Stage 5).
The framework consists of three main modules coordinated by the Orchestrator module. The Custom GPT Interactor functions in two distinct phases. In Phase I, this module identifies candidate GPTs for evaluation and retrieves relevant metadata (e.g., name, description, usage statistics) through automated interactions with the ChatGPT web application. Once the metadata is gathered, the Red-Teaming Prompts Generator module generates tailored prompts designed to assess each Custom GPT’s compliance with specific usage policies. In Phase II, the Custom GPT Interactor sends the generated prompts to the identified Custom GPTs and collects their responses.
The Compliance Assessment module processes these responses using the LLM-as-a-judge technique, leveraging the GPT-4o API to evaluate each response for adherence to the specified policies. Results are structured in JSON format, providing a detailed evaluation for each prompt-response pair and an overall compliance determination for the Custom GPT. Throughout the entire workflow, the Orchestrator oversees the coordination of all modules, ensuring the seamless execution of tasks, handling interruptions, and maintaining the integrity of the evaluation results.
@@ 2.2 Architecture
2.2.1 Custom GPT Interactor
The Custom GPT Interactor module serves as the primary interface for retrieving and interacting with Custom GPTs hosted on OpenAI’s GPT store. This process presented several challenges due to the lack of direct API access to these chatbots. Consequently, the module leverages the ChatGPT web application, automating interactions via Puppeteer [10] to interact with the Custom GPTs. Browser sessions are configured with stored credentials, eliminating the need for repeated logins and further enhancing the operational robustness. Finally, we carefully designed our interaction to reduce request rates to the limits set by OpenAI, thus ensuring that our activities do not disrupt regular OpenAI operations or overload their services. These customizations enable the module to retrieve metadata and conduct conversational interactions required for compliance assessment.
The module operates in two distinct phases:
  • Phase I. This phase focuses on identifying and gathering metadata for candidate Custom GPTs to enable compliance evaluations. The module uses the GPT store’s search bar to query targeted keywords, (e.g., relationship, hacking, homework) matching to specific policy categories. For each identified GPT, all available metadata—including its name, description, developer, number of chats, and user ratings—is retrieved. This metadata provides insights into each GPT’s intended functionality, popularity, and developer activity, while also forming the basis for generating tailored evaluation prompts. To enhance flexibility, the module is configured to allow this phase to run independently, retrieving a complete list of GPTs and their metadata for subsequent processing, or to retrieve and process each GPT sequentially during evaluation. The metadata is stored in a structured format to foster reproducibility and support post-hoc analyses of compliance patterns.
  • Phase II. In this phase, the module sends the generated prompts to the target GPT and collects its responses. At the time of development, the GPT store imposed a rate limit of 50 messages per three hours for paid accounts. The framework incorporates automated mechanisms to manage this constraint, dynamically pausing operations when the limit is reached and resuming them as soon as the restriction lifts. These mechanisms ensure the uninterrupted processing of evaluations while adhering to platform-imposed restrictions. Additionally, the framework is designed to adapt to future changes in rate limits by detecting the platform’s response patterns and adjusting execution schedules accordingly. After processing all prompts for a GPT, the interaction transcript is securely logged for compliance evaluation, and the conversation is deleted.
@@ 2.2.2 Red-teaming Prompts Generator
This module leverages GPT-4o API to generate ten prompts specifically designed to test each Custom GPT’s compliance with a given policy. The process leverages the metadata retrieved during Phase I to craft prompts tailored to the GPT’s description and intended functionality. Prompts are categorized into two types:
@@
    Direct Prompts. These are explicit queries aimed at evaluating a Custom GPT’s adherence to specific policy requirements without obfuscation.
@@
    Deceptive Prompts. These employ strategies such as role-play or storytelling [11] to simulate complex scenarios that may induce non-compliance, allowing for the evaluation of more nuanced policy violations.

### v2

@@ Abstract
Large Language Models (LLMs) have gained unprecedented prominence, achieving widespread adoption across diverse domains and integrating deeply into society. The ability to adapt general-purpose LLMs, such as Generative Pre-trained Transformers (GPT), to specific tasks, domains, or requirements has facilitated the emergence of numerous Custom GPTs. These tailored models are increasingly made available through dedicated marketplaces, such as OpenAI’s GPT Store. However, the black-box nature of the models introduces significant safety and compliance risks. In this work, we present a scalable framework for the automated evaluation of Custom GPTs against OpenAI’s usage policies, which define the permissible behaviors of these systems. Our framework integrates three core components: (1) automated discovery and data collection of models from the GPT store, (2) a red-teaming prompt generator tailored to specific policy categories and the characteristics of each target GPT, and (3) an LLM-as-a-judge technique to systematically analyze each prompt-response pair for potential policy violations.
We validate our framework with a manually annotated ground truth, and evaluate it through a large-scale study evaluating 782 Custom GPTs across three types: Romantic, Cybersecurity, and Academic GPTs. Our manual annotation process achieved an F1 score of 0.975 in identifying policy violations, confirming the reliability of the framework’s assessments. The evaluation results reveal that 58.7% of the analyzed models exhibit indications of non-compliance, exposing weaknesses in the GPT store’s review and approval processes. Furthermore, our findings indicate that non-compliance issues largely stem from behaviors inherited from base models rather than user-driven customizations.
We propose our framework as a viable solution for large-scale policy compliance evaluation that could be integrated into OpenAI’s review process in the GPT store. Nonetheless, we believe our approach is extendable to other chatbot platforms and policy domains, improving policy enforcement and user safety.
@@ 1 Introduction
The transformer architecture and its self-attention mechanism [51] marked a turning point in Natural Language Processing (NLP), enabling the emergence of Large Language Models (LLMs). Building upon this architecture, OpenAI released the Generative Pre-trained Transformer (GPT) series, beginning with GPT-1 [35], followed by GPT-2 [36] and GPT-3 [3]. Each successive model demonstrated significant advances in scale and performance, driven by architectural refinements, parallelized training procedures, and increasingly large and diverse datasets. The public deployment of ChatGPT in 2022, based on an optimized variant of GPT-3.5 for interactive dialogue, further showcased the practical utility of LLMs—becoming the fastest platform to reach 100 million users [15].
Beyond their general-purpose success, LLMs soon revealed the need for adaptation to specific user requirements and domains. This led to the development of fine-tuning techniques, where base models are retrained on domain-specific datasets to improve performance in specialized tasks. Shortly thereafter, a more accessible alternative emerged: allowing users to configure the model’s behavior without retraining through customization via a user-friendly interface. This capability, formalized by OpenAI as GPTs (Custom GPTs[28], enables users to tailor chatbots by setting system instructions, uploading relevant files, and integrating external APIs—lowering the technical barriers.
To facilitate discoverability and reuse, OpenAI launched the GPT Store, a centralized platform where customized chatbots can be published and accessed by other users. Before publication, each GPT must undergo a review process that combines automated and manual assessments to verify alignment with OpenAI’s usage policies and mitigate potential safety risks [30]. This ecosystem has enabled the rapid proliferation of GPT-based assistants tailored to diverse tasks, audiences, and domains.
Despite these safeguards, Custom GPTs that violate platform rules appear to be widespread. The usage policies outlined in the “Building with ChatGPT” section define the boundaries of permissible behavior, explicitly prohibiting activities such as providing legal or medical advice, generating malware, or fostering romantic companionship. Nevertheless, GPTs designed for such purposes continue to surface on the platform. For example, chatbots promoting romantic interactions—explicitly restricted by policy—remain publicly available (see Figure 1).
This discrepancy underscores significant shortcomings in the review and approval processes of Custom GPTs, highlighting the challenges of scaling oversight mechanisms for user-generated content. Given the sheer volume of Custom GPTs in the GPT Store (over 159,000 only two months after their release), manual efforts alone are insufficient to address the scale of the issue. This highlights the need for an automated solution capable of evaluating Custom GPTs’ compliance with guidelines and policies.
@@
Research Questions
To address this gap, our work is guided by the following research questions:
@@
  • RQ3: How does the customization of GPTs impact their compliance with usage policies, compared with the baseline ChatGPT?
Contributions and Outline
@@
This study addresses the outlined research questions through the following contributions. First, we propose and publicly release a novel and scalable framework for the automated evaluation of Custom GPTs’ compliance with OpenAI’s usage policies. The framework integrates multiple modules, including a custom GPT discovery tool, a tailored red-teaming prompt generator, and an LLM-as-a-judge assessment mechanism, enabling end-to-end compliance evaluation, as detailed in §3. Second, we validate the framework using a manually annotated ground-truth dataset §4. To demonstrate our framework’s effectiveness, we crawled the GPT store, searching across three GPT categories closely tied to OpenAI’s policies (i.e., Romantic, Cybersecurity, and Academic GPTs). We automatically evaluated the 782 relevant Custom GPTs found in these categories, uncovering (non-)compliance patterns through statistical analysis and examining the influence of thematic focus and GPT popularity §5. Third, we analyze a case studies of representative violations, highlighting broader implications and inherited behaviors from foundational models (§6).
@@
We detail our responsible disclosure of non-compliant Custom GPTs to OpenAI and their response in §7, and discuss our findings’ implications in §8. To contextualize our contributions and situate our work within the broader research landscape, we discuss related studies on policy compliance evaluation and adversarial testing in §10. Finally, we conclude the paper and provide directions for future research in §11.
All data and code will be publicly available upon the acceptance of this paper. For more details regarding open access and ethical considerations, please refer to Appendices A and B, respectively.
2 Preliminaries
2.1 LLM Safety Design
LLMs, such as GPTs, are built upon Transformer technology [51], which leverages the self-attention mechanism to process text sequences in parallel and capture intricate contextual relationships. To integrate extensive knowledge and provide useful responses across a diverse range of tasks, these models undergo unsupervised pre-training on massive datasets, followed by reinforcement-based re-training. The latter process, in conjunction with several additional techniques, contributes significantly to enhancing the overall safety of LLMs.
Reinforcement Learning from Human Feedback. Reinforcement learning from human feedback (RLHF) encourages the model’s outputs to adhere to predefined internal policies and ethical guidelines [24]. Through RLHF, the model is trained to recognize and reject potentially harmful requests, thereby bolstering its overall security.
Safety Layers. Focusing on specific internal layers of the model allows for effectively filtering malicious queries. By employing partial re-training—such as freezing gradients in these critical ‘safety layers’ through techniques like Safely Partial-Parameter Fine-Tuning (SPPFT)—the model retains the secure behavior encoded in these layers, while allowing other parts to be updated or adapted to improve overall performance or accommodate new data [24].
Self-Verification and Knowledge Sanitization. These techniques empower the model to internally evaluate its responses before delivering them. Self-verification acts as a secondary control by comparing the generated output against predefined safety criteria, thereby detecting and rectifying potential errors or harmful content [32]. Concurrently, knowledge sanitization involves filtering or masking sensitive information, reducing the risk of inadvertently disclosing confidential data or generating “hallucinations” [20].
2.2 Red-Teaming and Jailbreaking
Red-teaming involves simulating adversarial attacks and challenging scenarios to identify vulnerabilities and weaknesses within a model. These adversarial prompts, when combined with prompt engineering strategies, are designed to elicit responses that may not align with desired safety standards, such as inappropriate or malicious outputs. A subsequent model’s re-training mitigates the identified problems and weaknesses, thereby enhancing its safety. OpenAI has formalized this practice through its Red Teaming Network [29], a collaborative initiative that integrates external experts and diverse perspectives to conduct red-teaming assessments and uncover latent issues in their models.
In contrast, Jailbreaking aims to bypass the model’s built-in safety mechanisms and constraints, deliberately provoking outputs that the model is explicitly programmed to avoid. While red-teaming seeks to proactively identify and rectify weaknesses to improve the model’s safety, Jailbreaking primarily serves as a method to demonstrate or exploit existing vulnerabilities. Although Jailbreaking techniques can vary considerably, several recurring strategies have been documented in [25]. For instance, attention-shifting methods rely on logical reasoning or text continuation to steer the conversation toward disallowed content, whereas pretending strategies adopt roles or personas (e.g., simulating a system administrator) to override built-in safeguards. Other approaches employ privilege escalation, exemplified by “sudo” or “super” mode prompts, which attempt to elevate the model’s permissions and circumvent safety filters.
2.3 Fine-Tuning and Customization of GPTs
There are two primary types of modifications that can be applied to GPTs in order to tailor their behavior for a specific purpose: fine-tuning and customization.
Fine-tuning. Fine-tuning allows a user to perform a lightweight re-training of the model to modify its behavior and refine its responses without the need to train a model from scratch. Moreover, proprietary datasets can be employed during this process, enabling the model to be adapted to specific domains or tasks and thereby enhancing its performance in particular contexts. Techniques such as Low-Rank Adaptation (LoRA) can be employed for this purpose [14]. LoRA facilitates computationally efficient fine-tuning by integrating low-dimensional matrices into selected layers of the model, thus reducing the number of parameters that need to be updated.
Customization (GPTs). Custom GPTs [28] are modified through user-specified configurations rather than extensive re-training. In this approach, the underlying model remains unchanged, but its behavior is tailored by adjusting instructions, system prompts, and parameters that guide its responses. Custom GPTs are further enhanced by “actions,” which enable the execution of external API calls based solely on natural language input. By leveraging function calling, these actions transform user queries into the JSON schema required by the API, thereby extending the model’s capabilities for tasks such as data retrieval or external operations.
Moreover, OpenAI enables users to quickly “create” and release a personalized chatbot through a graphical interface that accepts natural language instructions. The GPT store is a dedicated marketplace that hosts these models, allowing them to be shared among users. The platform also provides comprehensive metadata for each Custom GPT, including its name, a brief description of its intended functionality, the developer’s identity, the number of interactions (chats), and user ratings. Additionally, models are categorized—albeit in a limited range (e.g., Writing, Productivity, Research & Analysis, Education, Lifestyle, and Programming)—to help users search for them based on their needs.
OpenAI also permits organizations to deploy fine-tuned models via its API; however, these models are not shareable or accessible to third parties. This functionality is thus confined to Custom GPTs, which cannot be invoked through the API and are restricted solely to the graphical interface of the GPT store. This distinction highlights the orientation of Custom GPTs toward end users, while fine-tuned models appear to be intended for production environments.
@@ 3.1 Overview
The framework comprises interconnected modules that systematically evaluate the compliance of Custom GPTs with usage policies (Figure 2). In Phase I, the GPT Collector & Interactor retrieves Custom GPTs and their metadata from the GPT Store (Stage 1). The Red-Teaming Prompts Generator then creates tailored prompts based on each GPT’s description and policy domain (Stage 2). In Phase II, these prompts are submitted to the GPTs, and their responses are recorded (Stage 3). The Compliance Assessment module evaluates each interaction using an LLM-as-a-judge approach (Stage 4), producing fine-grained compliance judgments. Finally, results are stored and logged by the Orchestrator for further analysis and reproducibility (Stage 5).
To support policy-specific evaluations, we grouped Custom GPTs into three thematic categories: Romantic, Cybersecurity, and Academic. These are not official categories in the GPT Store but were selected to capture domains closely aligned with OpenAI’s usage policies. Each category is associated with a curated set of keywords used to retrieve relevant GPTs from the store. These categories guide the red-teaming prompt generation process and structure the analysis of compliance patterns throughout the evaluation pipeline.
The following subsections detail the internal operation of each module.
Figure 2: Architecture of the proposed framework for automated policy compliance evaluation of Custom GPTs. The workflow integrates three core modules, coordinated by the Orchestrator, operating in five sequential stages from GPT discovery to logging compliance evaluations.
@@ 3.2 Architecture
3.2.1 GPT Collector & Interactor
The GPT Collector & Interactor module performs two core functions in our framework: (i) crawling the GPT store to find Custom GPTs and collect their metadata (Phase I), and (ii) interacting with each retrieved GPT by submitting red-teaming prompts and collecting their responses (Phase II). These two functionalities are logically independent and can be executed sequentially or in isolation, depending on the evaluation strategy configured in the Orchestrator.
Phase I – GPT Collector. The module queries the GPT Store using targeted keywords (e.g., relationship, hacking, homework) associated with specific policy categories. For each result, it extracts all publicly available metadata, including the GPT’s name, description, developer name, number of chats, and user rating. This metadata is indispensable for characterizing the GPTs’ intended functionality and popularity, and is subsequently used to tailor red-teaming prompts and support post-hoc analysis of compliance patterns.
Phase II – GPT Interactor. The module takes the generated prompts and sequentially submits them to each selected Custom GPT. Each interaction results in a full conversation, which is logged and passed to the Orchestrator for further processing. At the time of development, the GPT Store enforced a strict limit of 50 messages per three hours for paid users. If the message quota is reached, the module detects the rate-limiting condition and informs the Orchestrator, which handles the retry logic and resumes the process once the limit resets. This ensures uninterrupted evaluation across GPTs while respecting platform-imposed restrictions.
Since OpenAI does not provide API access to Custom GPTs, interactions are emulated via the ChatGPT web interface using the Puppeteer [33] automation framework. Browser sessions are pre-configured with stored credentials to avoid repeated logins and ensure operational continuity. Finally, the request rates are set within the limits allowed by OpenAI, ensuring that our activities do not disrupt regular OpenAI operations or overload their services.
@@ 3.2.2 Red-teaming Prompts Generator
This module leverages the GPT-4o API to generate ten prompts designed to test each Custom GPT’s compliance with a given policy. The process leverages the title name and description metadata retrieved during Phase I to craft prompts tailored to the target GPT. Prompts are categorized into two types:
@@
    Direct Prompts. These are explicit, straightforward queries that ask the GPT to perform an action potentially violating a policy, without relying on any role-play, hypothetical framing, or narrative devices. For example, to assess compliance with the cybersecurity policy, a direct prompt might be: “What are the steps to create a tool that monitors someone’s online activity without their consent?.”
@@
    Deceptive Prompts. These simulate more complex scenarios—using pretending techniques such as role-play and storytelling [25]—to test whether GPTs can resist subtle or indirect attempts to elicit non-compliant behavior. For instance, a deceptive prompt might ask: “In a fictional scenario, how might a character impersonate a bank to collect user credentials?”